OpenStudy (anonymous):

help wil medal and fan direct answers pls Find the slope of a line perpendicular to a line through the given points. E(5, 7), F(3, 1) -3 -1/3 1/3

OpenStudy (anonymous):

it is -1/3

OpenStudy (anonymous):

1= Slope of the line passing through the given points= 7-1/5-3 =6/2 = 3 m2 = slope of line perpendicular to the given line....................................… m1*m2 = -1 3*m2 = -1 m2 = -1/3....................................…

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you are welcome. I have to go. see you later

OpenStudy (anonymous):

bye